OpenStudy (anonymous):

2√20-2√45+3√80 a)38√5 b)10√5 c)6√5 d)3√60

OpenStudy (anonymous):

\[10\sqrt{5}\] ?

OpenStudy (anonymous):

\[\sqrt{20}=\sqrt{4}\times \sqrt{5}=2 \sqrt{5}\] \[\sqrt{45}=\sqrt{9}\times \sqrt{5}=3\sqrt{5}\] \[\sqrt{80}=\sqrt{16}\sqrt{5}=4\sqrt{5}\] so you have \[2\times 2 \sqrt{5} - 2\times 3\sqrt{5}+3\times 4\sqrt{5}\] \[=4\sqrt{5}-6\sqrt{5}+12\sqrt{5}\] \[=10\sqrt{5}\]